(1 point) Let A-0 -2 3 Find a basis of nullspace(A). Answer: To enter a basis into WeBWorK, place the entries of each vector inside of brackets, and enter a list of these vectors, separated by commas. For instance, if your basis is 21 , then you would enter [1,2,3],11,1,1] into the answer blank.
Answer #1

matrix is

0 4 6
0 -2 -3
0 6 9

convert into Reduced Row Eschelon Form...

Divide row1 by 4

0 1 3/2
0 -2 -3
0 6 9


Add (2 * row1) to row2

0 1 3/2
0 0 0
0 6 9


Add (-6 * row1) to row3

0 1 3/2
0 0 0
0 0 0

reduced matrix is

0 320 0 100

x=x.......................free variable

y+\frac{3}{2}z=0................y=-\frac{3}{2}z

z=z.......................free variable

.

solution is

\begin{pmatrix}x\\ y\\ z\end{pmatrix}=x\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}+y\begin{pmatrix}0\\ -\frac{3}{2}\\ 1\end{pmatrix}

null space are:

{\color{Red} \begin{pmatrix}1\\ 0\\ 0\end{pmatrix},\:\begin{pmatrix}0\\ -\frac{3}{2}\\ 1\end{pmatrix}}